OpenStudy (anonymous):

Which statements are true, if f(x) = x2 and g(x) = –x? Check all that apply. For any value of x, f(x) > g(x). For any value of x, f(x) < g(x). When x = –3, f(x) < g(x). When x = 2, f(x) > g(x). For all positive values of x, f(x) > g(x). For all negative values of x, f(x) > g(x).

OpenStudy (gorv):

yeah as f >g alwys not true as for x= -1/2 f<g

OpenStudy (gorv):

for x= 1 f>g so f<g not alwys true

OpenStudy (gorv):

at x=-3 f=(-3)^2=9 g=-(-3)=3 f>g not less than g

OpenStudy (gorv):

at x=2 f=4 g=-2 so f>g

OpenStudy (gorv):

For all positive values of x, f(x) > g(x).true as we seen above For all negative values of x, f(x) > g(x ) false for example x=-1/2 f=1/4 g=1/2 g>f